OpenStudy (anonymous):

solve system by substituion 6x-y=31 and 4x+3y=17

OpenStudy (amistre64):

6x - y = 31 4x+3y = 17 y = [17(6)-31(4)] / [3(6) - -1(4)] = (102-124) / (18+4) = -22 / 22 = -1 -1 +6x = 31 3y +4x = 17 x = [17(-1)-31(3)] / [4(-1)-6(3)] = (-17-93) / (-4-18) = 110 / 22 = 10/2 = 5